OpenStudy (kittiwitti1):

The displacement of a particle on a vibrating string is given by the equation s(t)=4+4sin(8πt), where s is measured in centimeters and t is in seconds. Find the velocity and acceleration of the particle after t seconds. I am aware that velocity involves a first derivative and acceleration involves the second derivative. LaTeX will be below.

iYuko (iyuko):

Well.....

OpenStudy (kittiwitti1):

First derivative:\[s'(t)=(4+4\sin{8\pi t})'\]\[=0+4(\cos{(8\pi t)}\times8\pi)\]\[=4(\cos{(8\pi t)}\times32\pi\]

OpenStudy (kittiwitti1):

Oops, left out a parentheses in the final part. How is the result? Is this correct?

OpenStudy (agent0smith):

You seem to have an extra 4 in there

OpenStudy (kittiwitti1):

I do? Where D:

OpenStudy (agent0smith):

Check your work in the second and third lines.

OpenStudy (kittiwitti1):

...I don't get it D:

OpenStudy (agent0smith):

You don't notice an 8 magically turning into a 32?

OpenStudy (kittiwitti1):

Well, yeah, it's \(4\times8\pi\) isn't it? o-o

OpenStudy (agent0smith):

Do you not see it? This is what you wrote earlier\[\large 4(\cos{(8\pi t)}\times8\pi)\] \[\large 4(\cos{(8\pi t)}\times32\pi)\]

OpenStudy (kittiwitti1):

I said I missed a parentheses in the last bit, but I meant it as \[4(\cos{(8\pi t)})\times32\pi\]

OpenStudy (agent0smith):

Math doesn't work that way.

OpenStudy (agent0smith):

\[\large a(b*c) \neq ab*ac\]

OpenStudy (kittiwitti1):

OH, wow. That was stupid of me.

OpenStudy (kittiwitti1):

\(s'(t)=32\pi\cos{(8\pi t)}\)?

OpenStudy (agent0smith):

Yes.

OpenStudy (kittiwitti1):

Alright, thank you! And the derivative of that....\[s"(t)=-256\pi^{2}\sin{(8\pi t)}\]Is this?

zepdrix (zepdrix):

LOL using quotation mark for second derivative, that's cute XD

OpenStudy (agent0smith):

^lol Looks about right. You can use wolfram alpha to find derivatives btw.

OpenStudy (kittiwitti1):

I believe I put two apostrophes :\ oh well. alright thanks for the tip!

zepdrix (zepdrix):

\(\large\rm "~\ne~~''\) XD

OpenStudy (kittiwitti1):

I guess it must've been my auto correct then. Oops.